john uses the accrual basis for his cleaning business. He received $2,400 on july 1, 2019 to perform cleaning for the 2 year period beginning on july 1, 2019. what amount of income must john report in:

2019__________

2020__________

2021__________
Answer #1

In 2019 = $2,400 X 6/24 = $600

In 2020 = $2,400 X 12/24 = $1,200

In 2021 = $2,400 X 6/24 = $600